OpenStudy (anonymous):

20x-3[2x+5(x-1)]-(x-1)

OpenStudy (anonymous):

28x-16

OpenStudy (anonymous):

20x-3(2x+5x-5)-x+1=20x-6x-15x+15-x+1=-2x+16=2(8-x)